A four-storey industrial building has a depreciated value of P890,000 at the start of 18th year and has a salvage value P50,000. What is the total depreciation during the 8th year? Use double declining balance method.


1
Expert's answer
2021-11-07T19:43:36-0500
"\frac{890000-50000}{18}=46667"

"\frac{46667}{890000}\times 100=5.2"

1st


"2\times0.052\times890000=92560"

2nd


"2\times0.052\times(890000-92560)=82834"

3d


"2\times0.052\times714606=74319"

4th


"2\times0.052\times640287=66590"

5th


"2\times0.052\times573697=59664"

6th


"2\times0.052\times514033=53459"

7n


"2\times0.052\times460574=47900"

8t


"2\times0.052\times412674=42918"